OpenStudy (anonymous):

Reduce each of the following to a function of a positive angle less than 45 degrees: a. sin 175 degrees

OpenStudy (anonymous):

@myininaya

myininaya (myininaya):

so are you saying we want to convert to sin(x) such that x is between 0 and 45 degrees?

OpenStudy (anonymous):

yep.. how dude :D

myininaya (myininaya):

http://www.mathsisfun.com/geometry/images/circle-unit-304560.gif Look at this unit circle... What do you notice about sin(150) and sin(30)?

OpenStudy (anonymous):

THE sign is different.

myininaya (myininaya):

Look again.. sin(150)=sin(30)=1/2

OpenStudy (anonymous):

oh.. yeah

myininaya (myininaya):

Both sin(150)=sin(30) and sin(180-150)=sin(30) ... what about sin(135) and sin(45) or sin(120) and sin(60)

myininaya (myininaya):

So we have sin(150)=sin(180-150)=sin(30) We also have sin(135)=sin(180-135)=sin(45) And sin(120)=sin(180-120)=sin(60) You should be able to use this pattern to make a guess what sin(175) equals

OpenStudy (anonymous):

how?

myininaya (myininaya):

All of those angles I just looked at or between 90 and 180... 175 is between 90 and 180

myininaya (myininaya):

the pattern should still follow

myininaya (myininaya):

Well I looked at angles from 90 to 180 and compared them to angles from 0 to 90

myininaya (myininaya):

But each time what did we do?

myininaya (myininaya):

You should a number that isn't changing in the process...

OpenStudy (anonymous):

so cos 5 ?

myininaya (myininaya):

sin(150)=sin(180-150)=sin(30) sin(135)=sin(180-135)=sin(45) sin(120)=sin(180-120)=sin(60)

OpenStudy (anonymous):

or i will use sin 5 ?

myininaya (myininaya):

Well is sin(150)=sin(30) or cos(30)?

myininaya (myininaya):

Look at the unit circle.

OpenStudy (anonymous):

sin 180 - 5

OpenStudy (anonymous):

it is sin

myininaya (myininaya):

sin( ?)

myininaya (myininaya):

What is your final answer?

OpenStudy (anonymous):

because i use 180 degrees? so its sin 5?

myininaya (myininaya):

yes sin(5)

OpenStudy (anonymous):

oh yeah i got it

OpenStudy (anonymous):

but how about cot 321degrees 54 min.

OpenStudy (anonymous):

so i will use sin360 - sin38.6 = -sin 38.6?